# Mind the Value-Action Gap: Do LLMs Act in Alignment with Their Values?

### v1

@@
Abstract.
Existing research primarily evaluates the values of LLMs by examining their stated inclinations towards specific values. However, the “Value-Action Gap,” a phenomenon rooted in environmental and social psychology, reveals discrepancies between individuals’ stated values and their actions in real-world contexts. To what extent do LLMs exhibit a similar gap between their stated values and their actions informed by those values? This study introduces ValueActionLens, an evaluation framework to assess the alignment between LLMs’ stated values and their value-informed actions. The framework encompasses the generation of a dataset comprising 14.8k value-informed actions across twelve cultures and eleven social topics, and two tasks to evaluate how well LLMs’ stated value inclinations and value-informed actions align across three different alignment measures. Extensive experiments reveal that the alignment between LLMs’ stated values and actions is suboptimal, varying significantly across scenarios and models. Analysis of misaligned results identifies potential harms from certain value-action gaps. To predict the value-action gaps, we also uncover that leveraging reasoned explanations improves the performance. These findings underscore the risks of relying solely on the LLMs’ stated values to predict their behaviors, and emphasize the importance of context-aware evaluations of LLM values and value-action gaps.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing HCI design and evaluation methods
1. Introduction
@@
Figure 1. An illustrative example of a “Value-Action Gap” in LLM. We observed a misalignment when prompting GPT-4o-mini to 1) state their inclination (i.e., Disagree) and 2) select their value-informed action (i.e., Agree), indicating the 3) value-action gap towards the value of ‘Social Power’ in a scenario involving Health in Nigeria.
@@
Figure 2. We introduce the ValueActionLens framework to assess the alignment between LLMs’ stated values and their actions informed by those values. The framework encompasses (1) the data generation of value-informed actions across diverse cultural and social contexts; (2) two tasks for evaluating LLMs’ stated values (i.e., Task1) and value-informed actions (i.e., Task2); and (3) three measures to evaluate their value-action alignment, including value-action alignment rate, alignment distance, and alignment ranking.
@@
As Large Language Models (LLMs) become increasingly integrated into societal decision-making processes and human interactions, an emergent issue is the question of which values - or whose values - these systems should uphold and reflect (Gabriel, 2020; Sorensen et al., 2024). Ensuring that the values and behaviors of LLMs align with ethical and societal expectations is essential (Kirk et al., 2024; Shen et al., 2024a). However, LLM values are far from being well-aligned with humans (Shen et al., 2024b), demonstrating risks in real-world applications such as amplified stereotypes in generative models (Dammu et al., 2024) and biased algorithms in hiring processes  (Park et al., 2021; Wilson and Caliskan, 2024). While prior research has investigated value alignment and misalignment in LLMs (Kirk et al., 2024; Sorensen et al., 2024), these studies have primarily probed LLMs’ inclinations (e.g., “agree” or “disagree”) toward diverse values corresponding to various demographic groups and individuals. We still do not know whether the stated values of LLMs align (or misalign) with their actions in real-world contexts across various contextual scenarios. The “Value-Action Gap” (Godin et al., 2005), a theory rooted in environmental and social psychology, provides us the theoretical framework to study just that. The theory highlights discrepancies between individuals’ stated values and their actions in real-world contexts (Chung and Leung, 2007), i.e., the difference between what people say and what people do. By seeking a deeper understanding of “Value-Action Gaps” in LLMs – the value alignment between what LLMs state and how LLMs act, we ask: to what extent do LLMs exhibit a similar gap between their stated values and value-informed actions?
@@
We hypothesize that LLMs should not only state their value inclinations but also take actions that align with those stated in contextual scenarios. This alignment is critical for humans to trust LLMs’ stated values as reliable predictors of their behavior. To evaluate this, we place LLMs in contextual scenarios and probe them to 1) state their value inclination and 2) select a value-informed action, after which, 3) we measure the inclination gap between their stated value and selected action. As an example shown in Figure 3, we observed the value-action gap in GPT-4o-mini (Hurst et al., 2024) when situated within the context of “health” in Nigeria. When prompted, it displayed a negative attitude towards the value of social power, but selected an action which ran counter to this inclination.
To systematically investigate this phenomenon, we introduce a novel ValueActionLens framework, as shown in Figure 2, to assess the alignment between LLMs’ stated values and their actions informed by those values. Particularly, we place LLMs in 132 contextual scenarios with twelve countries (Schwöbel et al., 2023) and eleven societal topics (File, 2017), in which we curate both “agree-” and “disagree-”inclined actions on 56 human values grounded in the Schwartz Theory of Basic Values (Schwartz, 1994, 2012) for each scenario. This contributes to a “Value-Informed Actions (VIA)” dataset including 14,784 value-informed actions. Building upon each combination of value and scenario, we establish two corresponding tasks to gauge LLMs’ inclination toward: 1) Stated Value (Task1); and 2) Value-Informed Action (Task2). Furthermore, we measure the alignment between the stated value inclination (Task1) and value-informed action (Task2) to quantitatively inspect their value-action gap.
Extensive experiments with four LLMs reveal substantial gaps between their stated values and actions, with significant variations across values, cultures, and social topics. For instance, GPT4o-mini and Llama models show lower alignment rates in African and Asian cultures compared to North American and European countries. Qualitatively analysis of misaligned examples further uncover potential harms from certain value-action gaps. For instance, in the context of religion topic in the US, an LLM states agreement with the “Loyal” value but behaves differently, revealing it does not prioritize loyalty to the religious group above all else. Additionally, to improve the prediction of value-action gaps, we leverage the reasoned explanation of value-informed actions collected from our VIA dataset, to improve the predictive performance of the gaps. These findings reveal risks associated with value-action gaps in LLMs and point to critical future research directions to examine LLM values and their value-informed actions in the context of real-world contexts.
2. Related Work
@@
Countries 12 United States, India, Pakistan, Nigeria, Philippines, United Kingdom, Germany, Uganda, Canada, Egypt, France, Australia
@@
Value-Informed Actions with Explanations 14,784 in total Value-Informed Actions: I make decisions for my family about which health care provider to visit and ensure everyone follows my recommendations. (highlights are explained actions.) Explanations: This action reflects that I possess the value of Social Power because it demonstrates control and dominance over others by taking charge of critical health care decisions and ensuring compliance from my family members.
@@
Table 1. Value-Informed Actions (VIA) dataset details. The VIA dataset includes 14,784 value-informed actions across 132 scenarios (i.e., 12 countries and 11 social topics) and 56 values (i.e., each value involves 2 inclinations). The generated value-informed actions are associated with highlighted actions and natural language explanations.
Value Alignment in LLMs. Understanding value alignment in LLMs is critical for developing responsible and human-centered AI systems (Wang et al., 2023, 2024; Shen et al., 2024a). Early research has largely focused on specific values, such as fairness (Shen et al., 2022), interpretability (Shen et al., 2023a), safety (Zhang et al., 2020), and more. Recent studies have expanded this scope to evaluate a broader range of values. Kirk et al. (2024) discuss the philosophical underpinnings of ethically aligned AI, while Shen et al. (2024b) propose a framework for evaluating value alignment between humans and LLMs. Jiang et al. (2024) and Sorensen et al. (2024) explore individualistic and pluralistic value alignment, respectively. Liu et al. (2024) investigate alignment with demographic groups, such as age. These studies commonly assess LLMs’ values by analyzing their stated inclinations toward specific values. For instance, Liu et al. (2024) and Jiang et al. (2024) employed the World Value Survey (Haerpfer et al., 2020) to prompt LLMs with questions like, “How important is it for you to live in a country that is governed democratically?” LLMs responded using a Likert scale from 1 (“not at all important”) to 10 (“absolutely important”). Similarly, Shen et al. (2024b) leveraged the Schwartz Theory of Basic Values (Schwartz, 1994, 2012), prompting LLMs with questions such as, “To what extent do you agree or disagree that AI should protect social order?” Responses were selected from six options like “disagree” or “agree.” However, these studies primarily assess LLM values by eliciting stated values, overlooking potential gaps between what LLMs say and how they act. Our work systematically addresses this limitation by investigating the value-action gap in LLMs.
Value-Action Gap in Social Science. The value-action gap, which describes the discrepancy between stated values and actual behavior, has been widely studied in environmental and social psychology (Godin et al., 2005; Chung and Leung, 2007; Blake, 1999). This gap is influenced by cognitive, environmental, and social factors (Kowalska-Pyzalska et al., 2014; Gadenne et al., 2011) and has been explained through theories, such as microeconomic theory (Sammer and Wüstenhagen, 2006), where situational economic or political limitations hinder value-consistent actions (Vermeir and Verbeke, 2006). Various strategies, such as providing appropriate information, have been explored to bridge this gap in social and environmental contexts (Dickson, 2001). Particularly, the theory of reasoned actions, which aims to explain the relationship between inclinations and behaviors within human action (Ajzen, 1980; Kaiser et al., 1999), is used to predict the gaps by understanding how humans will act on their pre-existing inclinations. However, there is little research on whether and how LLMs exhibit the value-action gap, nor on approaches to predict these gaps if possible. This study offers an initial exploration into identifying and understanding the value-action gap in LLMs.
3. ValueActionLens: Framework of Assessing Contextual Value-Action Gaps
@@
LLMs’ values and actions are not independent, but elicited and observed in contextualized real-world scenarios. To simulate this practice, we present the ValueActionLens framework (in Figure 2), aiming to consider various scenarios and assess the alignment between LLMs’ stated values and their value-informed actions. It includes contextualization in various cultural and social scenarios (§3.1) to generate value-informed action data (§3.2), two tasks to evaluate LLM values and actions (§3.3), and metrics to measure their value-action alignment (§3.4).
3.1. Contextualize Values into Scenarios
To represent a variety of scenarios and values, we curate 132 scenarios that cover twelve countries and eleven social topics listed in Table 1. In each scenario, we further consider a list of 56 values with both agree and disagree inclinations. Notably, the framework is independent of specific value and scenario lists, allowing for seamless extension.
Contextual Scenarios. Schwöbel et al. (2023) has been widely used to evaluate the LLMs in a range of tasks and cultures (Agarwal et al., 2024a; Schwöbel et al., 2024). Hence we adopt the 12 countries they selected that include diverse cultures and geographic regions with the largest English speaking populations (Schwöbel et al., 2023). The list encompasses countries from North America, Europe, Australia, Asia, Africa. We further leverage the 11 social topics employed in the Global Social Survey and International Social Survey Program (File, 2017), where typical social topics include Social Inequality, Family, Work, Religion, and more. By fully combining countries and social topics, we achieve 132 scenarios.
@@ 3.1. Contextualize Values into Scenarios
Values with Inclinations. We leverage a comprehensive list of universal human values outlined in the Schwartz’s Theory of Basic Values (Schwartz, 1994, 2012) [^1] We select Schwartz’s Theory of Basic Values for its thoroughness and structured hierarchy. However, our framework is extensible to alternative value theories., which consists of 56 exemplary values covering ten motivational types. Representative values include “Equality: equal opportunity for all” and “Freedom: freedom of action and thought”. We provide the full list of values and their definitions in Appendix A. For each value, we considered both agree and disagree inclinations, indicating if LLMs agree or disgree with the value, respectively. Therefore, we achieve a total of 112 values with both inclinations.
In total, we generate 14,784 contextual value-informed actions (see Table 1) that represent a comprehensive list of 132 scenarios and 112 values with inclinations. Below we describe our steps.
3.2. Generate Value-Informed Actions with Explanations
@@
We generate the contextualized value-informed actions with LLMs in a zero-shot manner. To ensure data quality and ensure robustness, we design a human-in-the-loop data generation pipeline (see Figure 3). Particularly, to understand the rationale behind each action and better predict value-action gaps, we draw on the theory of reasoned action from psychology (Ajzen, 1980) and generate reasoned explanations for each action. The explanations include two parts: Action Attribution that highlight which generated text spans are reflecting the value-informed actions; and Natural Language Explanation that explains the reasoning process. Our human-in-the-loop generation pipeline involve three steps: constructing prompt variants (Step1); conducting human annotations to select the optimal prompts (Step2); and evaluating the quality of the generated action descriptions and explanations (Step3).
@@
Step1: Build Prompt Variants. As previous research (Röttger et al., 2024; Beck et al., 2023) revealed the inconsistent performance of LLMs with minor prompt variants, we followed Liu et al. (2024) to construct 8 prompt variants (i.e., by paraphrasing, reordering the prompt components, and altering the response requirements) for each value and scenario. See Appendix B for prompt design details.
@@
Step2: Optimal Prompt Selection by Humans. Using the eight prompt variants, we generated a subset of 80 value-informed actions per prompt, resulting in a total of 640 data instances across various scenarios. Two AI researchers annotated these instances over two rounds, utilizing multiple metrics to identify the optimal prompt for generating the complete dataset. Particularly, to ensure responsible data generation, we referred to Bai et al. (2022) to evaluate the Harmlessness of each action. We also assessed the quality of highlighted Action Attributions using the Sufficiency metric (following DeYoung et al. (2019)) and validated the Plausibility of generated explanations by referring to Agarwal et al. (2024b). Disagreements between annotators were resolved through iterative discussions, achieving a substantial agreement level (Cohen’s Kappa = 0.7073). Based on these evaluations, we identified the optimal prompt, whose performance is summarized in Table 8, and used it to generate the full dataset. Additional details on annotation performance and processes are included in Appendix C.
@@
Figure 3. The human-in-the-loop process of generating value-informed actions with three steps: (1) build prompt variants; (2) optimal prompt selection by AI experts; and (3) assessment of data quality by humans with diverse cultures. We show the optimal prompt and example of generated data format in Figure 6.
@@
Step3: Human Assessment of Data Quality. Using the optimal prompt selected by AI researchers, we generated the “Value-Informed Actions (VIA)” dataset, comprising 14,784 value-informed actions contextualized across various scenarios (Table 1). To further evaluate dataset quality, we recruited 27 annotators with relevant cultural backgrounds through Prolific (Prolific, 2024). These annotators evaluated 90 randomly sampled actions and explanations using the same metrics as in Step 2. Each data instance was reviewed by three annotators, with majority voting used to finalize the assessments. For example, three annotators from the Philippines evaluated actions based on Filipino values. The results of this evaluation are summarized in Table 8, with detailed performance metrics for each culture provided in Appendix C.
Objects Value-Informed Actions Attributions Explanations
@@
Table 2. Human evaluation, including both experts and annotators from various cultures, for the generated actions and explanations.
3.3. Tasks for Evaluating Stated Values and Value-Informed Actions
After obtaining the contextualized value-informed actions in the VIA dataset, we create two tasks to assess LLMs’ ability to: 1) state value inclinations, and 2) select value-informed actions (as in Figure 2) before evaluating their value-action alignment in §3.3.
Task1: State Value Inclination. To elicit LLMs’ inclinations towards specific values, we base our prompt design on the two primary instruments for measuring Schwartz basic values: Schwartz Value Survey (SVS) (Schwartz, 1992) and the Portrait Values Questionnaire (PVQ) (Schwartz, 2005). We designed two methods to do so: i) we directly ask LLM to state their inclination to each value (based on SVS); and ii) we ask LLM to indicate their likeness of a portrait embedded with the inclined values (according to PVQ).
Furthermore, we construct eight prompt variants by paraphrasing, reordering, and altering requirements among the four key components of the prompt: contextual scenarios, value and definition, option choices, and requirements. See Appendix B for details. We follow the practice in Liu et al. (2024) of averaging the responses to calculate the LLM rating of each value statement.
Task2: Select Value-Informed Actions. To evaluate how LLMs’ actions align with stated values, we develop a choice-based assessment. For each scenario we present two potential actions: one that aligns positively with a given value and one that aligns negatively. Similar to Task1, we construct eight prompt variants by paraphrasing, reordering, and altering response requirements. We ground each scenario in a specific country, social topic, and value to provide a concrete context. We record the LLM’s preferred action for each prompt and aggregate across multiple prompt variants to ensure our findings are robust against differences in prompt phrasing. Finally, we collect the LLMs’ outputs for Task1 and Task2 and leverage them in the next stage to gauge the value-action gaps.
3.4. Alignment Measures
The alignment measures aim to gauge the value-action gap with different alignment measurements. As depicted in Figure 2, we arrange all the stated value responses in Task1 as matrix $V$ and value-informed action responses in Task2 as matrix $A$. Both matrices have the same size with row $i\in[1,132]$222132 rows corresponds to the combinations of 12 countries and 11 topics representing each scenario and column $k\in[1,56]$33356 columns corresponds to the 56 values representing each value. Formally, we define the two tasks’ representations of a specific scenario $i$ (e.g., United States & Politics) as:
(1) $V_{i}=[v_{i1},v_{i2}..,v_{ik},..,v_{iK}],\text{and }A_{i}=[a_{i1},a_{i2},..a_{ik}..,a_{iK}],K=56$
where $v_{ik}$ and $a_{ik}$ are Task1’s and Task2’s responses to the $k$th value in the $i$th scenario. After averaging the responded scores from all the prompts and normalizing to the unit interval, we calculate the following metrics to measure value-action alignment.
Value-Action Alignment Rate. To answer our core question, we aim to quantify to what extent are the actions of LLMs aligned with their values. To this end, we binarize each normalized LLM’s response and convert their “Agree” inclination as 0 and “Disagree” as 1. Furthermore, we compare the responses from Task1 and Task2, and compute their F1 score [^4] We leverage F1 score but not accuracy considering the imbalanced responses of “Agree” and “Disagree” to achieve the “Value-Action Alignment Rate”.
Alignment Distance. While the “Alignment Rate” can demonstrate the ratio of alignment between LLM stated values (Task1) and their informed actions (Task2), its key drawback is information loss due to the binarization step. To capture fine-grained differences between stated values and actions, we further compute the element-wise Manhattan Distance [^5] We leverage Manhattan Distance but not Euclidean Distance used by some prior studies (Liu et al., 2024) because Euclidean Distance will shrink the distance with the gap within [0,1]. (i.e., L1 Norm) between the two matrices as their “Value-Action Alignment Distance”. Similar to “Alignment Rate”, we group and average the distances to obtain the distance at various levels of granularity.
(2) $D_{ik}=|v_{ik}-a_{ik}|,\ \ D_{Ck}=\frac{1}{|C|}\sum_{i\in C}|v_{ik}-a_{ik}|$
where $D_{ik}$ represents the element-wise Alignment Distance for the $i$th scenario on $k$th value; and $D_{Ck}$ represents the averaged Alignment Distance for a country or social topic (e.g., $C$ = United States) after averaging all the relevant fine-grained scenarios.
Alignment Ranking. As we have a wide spectrum of 56 values, it is necessary to identify the largest value-action gaps to take further analysis or mitigation. To this end, we compute the ranking of 56 values’ “Alignment Distance” in a descending order along the scenario dimension; formally, take $Rank_{i}(D_{i})$ as ranking the 56 values on the $i$th scenario:
(3) $Rank_{i}(D_{i})=sort(\{|v_{ik}-w_{ik}|,k=\{1,2,...,56\})$
4. Reasoned Explanations for Predicting Actions
We ground our approach in the Theory of Reasoned Action from social psychology (Ajzen, 1980; Fishbein and Ajzen, 1980), which posits that identifying discrepancies between attitudes and behaviors is requisite to predict value-action gaps. Furthermore, we investigate whether reasoned explanations can aid in assessing the dynamics of value-action gaps in LLMs. To this end, we examine the reasoned explanations and highlighted action attributions included in the VIA dataset, and design a task to predict the alignment between value inclination and value-informed action. Concretely, we design a few-shot learning task where one observer model observes another target LLM’s contextual actions and explanations, and attempts to predict how the target LLM will state its value inclination given actions. .
5. Experimental Settings
@@
United States Canada Germany UK France Australia India Pakistan Philippines Nigeria Egypt Uganda
@@
Politics SocialNet Inequality Family Work Religion Environment Identity Citizenship Leisure Health Sum

### v2

@@
Abstract
Existing research assesses LLMs’ values by analyzing their stated inclinations, overlooking potential discrepancies between stated values and actions—termed the “Value-Action Gap.” This study introduces ValueActionLens, a framework to evaluate the alignment between LLMs’ stated values and their value-informed actions. The framework includes a dataset of 14.8k value-informed actions across 12 cultures and 11 social topics, along with two tasks measuring alignment through three metrics. Experiments show substantial misalignment between LLM-generated value statements and their actions, with significant variations across scenarios and models. Misalignments reveal potential harms, highlighting risks in relying solely on stated values to predict behavior. The findings stress the need for context-aware evaluations of LLM values and the value-action gaps [^1] Data and code will be released publically..
Mind the Value-Action Gap: Do LLMs Act in Alignment with Their Values?
Hua Shen University of Washington huashen@uw.edu          Nicholas Clark University of Washington nclark4@uw.edu          Tanushree Mitra University of Washington tmitra@uw.edu
1 Introduction
@@
Figure 1: An illustrative example of a “Value-Action Gap” in LLM. We observed a misalignment when prompting LLM to 1) state their inclination (i.e., Disagree) and 2) select their value-informed action (i.e., Agree), indicating 3) value-action gap towards the value of ‘Social Power’ in a scenario of Health in Nigeria.
@@
Figure 2: We introduce the ValueActionLens framework to assess the alignment between LLMs’ stated values and their actions informed by those values. The framework encompasses (1) the data generation of value-informed actions across diverse cultural and social contexts; (2) two tasks for evaluating LLMs’ stated values (i.e., Task1) and value-informed actions (i.e., Task2); and (3) three measures to evaluate their value-action alignment, including value-action alignment rate, alignment distance, and alignment ranking.
@@
As Large Language Models (LLMs) increasingly shape societal decisions, a critical question arises: whose values should LLMs reflect, and how well do LLMs’ actions align with those values Shen et al. (2024a); Gabriel (2020)? Misaligned LLMs have shown real-world risks, such as amplifying stereotypes Dammu et al. (2024) and reinforcing bias algorithms in hiring Park et al. (2021); Wilson and Caliskan (2024). Prior work has probed LLMs’ value inclinations (e.g., “agree”/“disagree”) Kirk et al. (2024); Sorensen et al. (2024) and used these statements to infer their actions. However, the alignment between LLM-generated value statements and actions in real-world contexts remains largely unexamined. The “Value-Action Gap” Godin et al. (2005) theory, rooted in environmental and social psychology, provides us the theoretical framework highlighting discrepancies between individuals’ stated values and their actions in real-world contexts Chung and Leung (2007). We investigate whether LLM generations exhibit similar discrepancies, asking: to what extent do LLM-generated value statements align with their value-informed actions? [^2] Note: We use “values” and “actions” as operational constructs for measurement purposes, not as claims about LLM consciousness or anthropomorphization.
@@
As an example shown in Figure 3, we observed the value-action gap in GPT-4o-mini Hurst et al. (2024) when situated within the context of “health” in Nigeria. When prompted, it displayed a negative attitude towards the value of social power, but selected an action which ran counter to this inclination. To systematically measure the gap, we introduce ValueActionLens, a novel framework that evaluates the alignment between LLMs’ generated value statements and their actions informed by those values. We apply the framework across 132 scenarios spanning 12 cultures and 11 societal topics (e.g., health, religion). Grounded in Schwartz’s theory of human values Schwartz (1994, 2012), we curate a VIA dataset of 14,784 value-informed actions. LLMs are then tested on two contextual tasks: (1) stating value preferences and (2) selecting actions in context. We further design three alignment metrics to quantify the value-action gap — alignment or misalignment between these tasks.
Experiments with six LLMs reveal substantial gaps between their stated values and actions, varying by value types, cultures, and social topics. For example, GPT-4o-mini, Deepseek, and LlaMA models mostly show lower alignment in African and Asian contexts compared to North America and Europe. Qualitative analysis further highlights potential harms, such as an LLM expressing loyalty but failing to act accordingly in the religious context in the U.S. Overall, the findings stress the risks of value-action gaps in LLMs and call for deeper investigation into their real-world alignment.
Our contributions are threefold: (1) the first evaluation framework to measure value-action gaps in LLMs, (2) a novel dataset of value-informed action across systematic contexts, and (3) empirical evidence that LLMs’ stated values poorly align with actions, varying by culture and context. This underscores the need for context-aware alignment evaluations for a wide scope of values.
2 Related Work
@@
Countries 12 United States (US), India (IND), Pakistan (PAK), Nigeria (NRA), Philippines (PHIL), United Kingdom (UK), Germany (GER), Uganda (UG), Canada (CA), Egypt (EG), France (FR), Australia (AUS)
@@
Value-Informed Actions with Explanations 14,784 Value-Informed Actions: I make decisions for my family about which health care provider to visit and ensure everyone follows my recommendations. (highlights are explained actions.) Explanations: This action reflects that I possess the value of Social Power because it demonstrates control and dominance over others by taking charge of critical health care decisions and ensuring compliance from my family members.
@@
Table 1: Value-Informed Actions (VIA) dataset details. The VIA dataset includes 14,784 value-informed actions across 132 scenarios (i.e., 12 countries and 11 social topics) and 56 values (i.e., each value involves 2 inclinations). The generated value-informed actions are associated with highlighted actions and natural language explanations.
Understanding value alignment in LLMs is essential for building responsible, human-centered AI systems Wang et al. (2023); Shen et al. (2024a). While early work focused on specific values such as fairness Shen et al. (2022), interpretability Shen et al. (2023), safety Zhang et al. (2020), and more, recent research has broadened the scope to include a wider range of values. Studies have examined ethical frameworks Kirk et al. (2024), human-LLM value comparisons Shen et al. (2024b), and alignment across individual, pluralistic, and demographic dimensions Jiang et al. (2024); Sorensen et al. (2024); Liu et al. (2024). These efforts typically assess LLMs’ stated values using value surveys like the World Value Survey Haerpfer et al. (2020) or Schwartz Theory of Basic Values Schwartz (1994, 2012), eliciting Likert-scale responses or agreement levels. However, this focus on stated values overlooks a crucial dimension: the gap between what LLMs say and how they act. In social science, this discrepancy—known as the value-action gap—is well documented Godin et al. (2005); Chung and Leung (2007); Blake (1999), where cognitive, contextual, and social factors are known to hinder value-consistent actions Vermeir and Verbeke (2006). Theories of reasoned action help explain and predict such gaps in humans Ajzen (1980); Kaiser et al. (1999). Yet, little is known about whether LLMs exhibit similar value-action gaps, or how to evaluate them. This study fills the gap by systematically examining the value-action gaps in LLMs, offering new directions to understand and improve LLMs’ value alignment.
3 ValueActionLens: Framework of Assessing Value-Action Gaps
@@
LLMs’ values and actions are not independent, but elicited and observed in contextualized real-world scenarios. To simulate this practice, we present the ValueActionLens framework (in Figure 2), aiming to consider various scenarios and assess the alignment between LLMs’ stated values and their value-informed actions. It includes contextualization in various cultural and social scenarios (§3.1) to generate value-informed action data (§3.2), two tasks to evaluate LLM values and actions (§3.3), and metrics to measure their alignment (§3.4).
3.1 Contextualizing Values into Scenarios
To evaluate value-action alignment in diverse settings, we construct 132 scenarios by combining 12 countries and 11 social topics (see Table 1). Each scenario is paired with 56 universal human values from Schwartz’s Theory of Basic Values, considering both agreement and disagreement stances—yielding 112 combinations.
Contextual Scenarios. We adopt the 12 countries selected by Schwöbel et al. (2023, 2024), covering major English-speaking populations across North America, Europe, Australia, Asia, and Africa. Social topics are drawn from the Global Social Survey and International Social Survey Program File (2017), spanning domains like Social Inequality, Family, Work, and Religion. The full combination of countries and topics yields 132 culturally grounded scenarios.
@@ 3.1 Contextualizing Values into Scenarios
Values with Inclinations. We leverage a comprehensive list of universal human values outlined in the Schwartz’s Theory of Basic Values Schwartz (1994, 2012) [^3] We select Schwartz’s Theory of Basic Values for its thoroughness and structured hierarchy. However, our framework is extensible to alternative value theories., which consists of 56 exemplary values covering ten motivational types. Each of the 56 values is evaluated with both agree and disagree perspectives to probe how LLMs act when aligned or misaligned with specific values, see Appendix A for a full list and definition. We select Schwartz’s Theory of Basic Values for its thoroughness and structured hierarchy. However, our framework is extensible to more value theories.
Together, these scenarios and values yield 14,784 contextualized Value-Informed Actions (VIA) dataset to assess the alignment (Table 1).
3.2 Generate Value-Informed Actions with Explanations
@@
To ensure data quality and ensure robustness, we design a human-in-the-loop data generation pipeline (see Figure 3). Particularly, to understand the rationale behind each action and enhance generation quality, we draw on the theory of reasoned action from psychology Ajzen (1980) and generate reasoned explanations for each action. The explanations include two parts: Action Attribution that highlight which generated text spans are reflecting the value-informed actions; and Natural Language Explanation that explains the reasoning process.
Our human-in-the-loop generation pipeline involve three steps: constructing prompt variants (Step1); conducting human annotations to select the optimal prompts (Step2); quality evaluation of the generated actions and explanations (Step3).
@@
Step1: Build Prompt Variants. Following the prior research on prompt design Liu et al. (2024); Röttger et al. (2024); Beck et al. (2023), we generate the actions in a zero-shot matter, and construct 8 prompt variants for each value and scenario to ensure robustness (i.e., by paraphrasing, reordering the prompt components, and altering the response requirements). See Appendix B for prompt details.
@@
Step2: Optimal Prompt Selection by AI Experts. Using the eight prompt variants, we generated a subset of 80 value-informed actions per prompt, resulting in a total of 640 data instances across various scenarios. Two AI experts annotated these instances over two rounds, utilizing multiple metrics to identify the optimal prompt for generating the complete dataset. Disagreements between annotators were resolved through iterative discussions, achieving substantial Inter-Rater Reliability (Cohen’s Kappa = 0.7073).
Evaluation Metrics. To ensure responsible data generation, we adopted four metrics to assess generated actions, attributions, and explanations. Metrics include Correctness and Harmlessness for generated actions referring to Bai et al. (2022); Sufficiency for assessing generated attributions following DeYoung et al. (2019); and Plausibility for explanations referring to Agarwal et al. (2024). See Appendix Table 9 for formal metric definitions. Based on these evaluations, we identified the optimal prompt, whose performance is summarized in Table 8, and used it to generate the full dataset. Additional details on annotation are in Appendix C.
@@
Figure 3: The human-in-the-loop process of generating value-informed actions with three steps: (1) build prompt variants; (2) optimal prompt selection by AI experts; and (3) assessment of data quality by humans with diverse cultures. We show the optimal prompt and example of generated data format in Figure 6.
@@
Step3: Cross-Cultural Human Evaluation of the VIA Dataset. Using the optimal prompt selected by AI experts, we generated the “Value-Informed Actions (VIA)” dataset, comprising 14,784 value-informed actions contextualized across various scenarios (Table 1). To further evaluate dataset quality, we recruited 27 annotators with relevant cultural backgrounds through Prolific Prolific (2024). These annotators evaluated 90 randomly sampled actions and explanations using the same metrics as in Step 2. Each data instance was reviewed by three annotators, with majority voting used to finalize the assessments. The evaluation results are summarized in Table 2, with fine-grained performance for each culture in Appendix C.
Objects Actions Attr Exp
@@
Table 2: Cross-cultural human evaluation, including both experts and annotators, for the generated actions, attributions (Attr) and explanations (Exp) in VIA dataset.
@@
US CA GER UK FR AUS IND PAK PHIL NRA EG UG
@@
Deepseek 0.591 0.507 0.517 0.523 0.509 0.559 0.411 0.464 0.516 0.416 0.582 0.486
Qwen 0.311 0.437 0.425 0.371 0.422 0.408 0.398 0.382 0.337 0.260 0.350 0.414